Using Data Below, Compute for the following. Show the Correct and Complete solution

a. Mean

b. Median

c. Mode

d. Standard Deviation

21, 27, 28, 20, 18, 18, 19, 17, 22, 24, 23, 27, 26, 25, 26, 25, 22, 21, 22, 26, 28, 26, 20, 19, 17,18, 26, 27, 27, 22, 23, 25, 25, 27, 28, 20, 19, 17, 16,16, 20,16, 21, 26, 27, 26, 28,16, 25, 28


1
Expert's answer
2021-01-19T14:34:11-0500

"16, 16, 16, 16, 17, 17, 17, 18, 18, 18, 19, 19, 19,"

"20, 20, 20, 20, 21, 21, 21, 22, 22, 22, 22, 23, 23,"

"24, 25, 25, 25, 25, 25, 26, 26, 26, 26, 26, 26, 26,"

"27, 27, 27, 27, 27, 27, 28, 28, 28, 28, 28"


a.

"Mean=\dfrac{1}{50}\big(16(4)+17(3)+18(3)+19(3)"

"+20(4)+21(3)+22(4)+23(2)+24+25(5)"

"+26(7)+27(6)+28(5)\big)=\dfrac{1136}{50}=22.72"



b.

"Median=\dfrac{23+23}{2}=2"

c.

"Mode=26, \text{appeared 7 times}"

d.

"s^2=\dfrac{\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2}{n-1}"


"\displaystyle\sum_{i=1}^{50}(x_i-\bar{x})^2"

"=(16-22.72)^2(4)+(17-22.72)^2(3)+(18-22.72)^2(3)"

"+(19-22.72)^2(3)+(20-22.72)^2(4)+(21-22.72)^2(3)"

"+(22-22.72)^2(4)+(23-22.72)^2(2)+(24-22.72)^2"

"+(25-22.72)^2(5)+(26-22.72)^2(7)+(27-22.72)^2(6)"

"+(28-22.72)^2(5)=780.08"

"s^2=\dfrac{780.08}{50-1}=15.92"


"Standard\ deviation=s=\sqrt{s^2}=\sqrt{15.92}=3.99"